# Training-and-prompt-free General Painterly Harmonization Using Image-wise Attention Sharing

### v1

@@ 4.1. Share-Attention Module
where $\odot$ denotes the Hadamard product. Afterward, Eq.(5) outputs the batch attention $\hat{A}\in\mathbb{R}^{(3\cdot m)\times d}$ containing the intact $A^{f}$, $A^{b}$, and $A^{comp}$ guided by the features of $z^{\text{b}}$ and $z^{\text{b}}$. The specially designed $\hat{M}$ can be visualized as:
@@
Here, $J\in\mathbb{R}^{m\times m}$ is an all-one matrix, and $\nu=-\infty$ minimizes the similarity between $Q$ and $K$ on the corresponding entry, keeping $A^{f}$, $A^{b}$ intact. It is worth noting that when setting $\alpha=-\infty$, $\beta=-\infty$, and $\gamma=1$, each row in Eq.(5) is equivalent to Eq. (4) as $Q^{comp}$, $Q^{f}$, and $Q^{b}$ can only attend to its counterpart $K^{comp}$, $K^{f}$, and $K^{b}$ without information from other images. Therefore, our proposed shared attention module can be viewed as an expansion of the self-attention layer with adjustable entries controlling image-wise attention sharing, which is also capable to be incorporated into other image editing methods (see Appendix  H for more details).
@@
Figure 3. Comparisons of different attention strategy with corresponding similarity mask (read with Fig. 2).
4.2. Similarity Reweighting
A default approach to allow $z^{\text{comp}}$ to attend to all three images simultaneously is to set $\alpha$, $\beta$, and $\gamma$ in $M$ to 1 with the expectation that $z^{\text{comp}}$ would be able to incorporate information from $z^{\text{b}}$ and $z^{\text{b}}$, achieving stylized and harmonized output. However, as illustrated in Fig. 3(a), the output image $I^{o}$ appears identical to the input composite image $I^{\text{comp}}$ because the $Q^{\text{comp}}$-$K^{\text{comp}}$ pairs exhibit overwhelming similarity as compared to other pairs, thus dominating the entire reconstruction process. To fuse the images more naturally, it is neccesary to disentangle the overwhelming similarity caused by the $Q^{\text{comp}}$-$K^{\text{comp}}$ pairs while maintaining its original semantic information.
@@ 4.2. Similarity Reweighting
To eliminate the overwhelming issue, we propose to remove the similarity calculation of ($Q^{\text{comp}}$, $K^{\text{comp}}$) by setting $\gamma=-\infty$. Forcing $Q^{\text{comp}}$ only be able to utilize the key and value ”shared” by others ($z^{\text{b}}$ and $z^{\text{b}}$), the result can be found in Fig. 3(b). The intriguing fact is that the output image only changes slightly even when the pasted ”corgi” in $I^{\text{comp}}$ has an absolutely different scale of background compared to the ”corgi” in $I^{f}$. We infer that the failure of Fig. 3(b) is that the pretrained diffusion model is robust enough to capture high-level semantic and structural information even with minor disturbances, such as differences in scale and position. As such, the self-attention layer can withstand perturbations, generating results that are similar to the original input.
To decipher which perturbation level can break the self-attention robustness but still generate high-quality results. A simple idea with attention injection has been widely adopted by previous research (Gu et al., 2024; Hertz et al., 2023; Lu et al., 2023b; Tumanyan et al., 2023), which can be viewed as a strong perturbation of self-attention through additional information from other images. However, directly injecting additional attention does not blend well with the background image. The reason behind the malfunction of attention injection toward the self-attention layer is that the foreground and the background images lie in different distributions, causing its attention similarity low on others and overwhelmed by themselves. Hence, our share attention module cannot force the foreground to adapt the background style, as shown in Fig. 3(b).
To mitigate the gap between two distinct image distributions, we assign different values to $\alpha$ and $\beta$, ensuring that the merging of features occurs with variant weights, which forces the model to guide the image reconstruction in a predetermined direction. Specifically, $\alpha$ should be set to be greater than $\beta$, urging our model places greater emphasis on stylistic background features of $z^{\text{b}}$ while maintaining the integrity of the foreground content. This methodology can be interpreted as an adaptation of the attention mechanism, achieved by altering the similarity relationship between features, a technique we call ”soft injection.” Unlike direct feature integration methods that rely on a copy-and-paste paradigm, soft injection provides weighted blending. Our observations reveal that soft injection facilitates superior image-wise self-attention editing, especially when compared with conventional prompt-wise cross-attention editing techniques (Gu et al., 2024; Hertz et al., 2023; Lu et al., 2023b; Tumanyan et al., 2023), thus giving more granular control over the editing process. We provide an illustration in Fig. 3(c) by setting $\alpha=0.9$ and $\beta=1.1$, we can strike a satisfactory balance, harmoniously blending the foreground and background features. The overall algorithm can be found in Appendix  A.
@@
Figure 4. Qualitative result of object insertion (rows 1 and 2), object swapping (rows 3 and 4), and style transfer (rows 5-7)
@@
Table 1. Quantitative results of GPH-Benchmark ( represents the method with inference-time-adjustable hyperparameters. The left side of / represents content emphasis strategy, while the right side of / represents stylized emphasis strategy.
Painterly Harmonization Style Transfer
Ours TF-ICON(Lu et al., 2023b) PHDiff(Lu et al., 2023a) PHDNet(Cao et al., 2023a) SDEdit(Meng et al., 2022) Ours QuantArt(Huang et al., 2023) VCT(Cheng et al., 2023) StyTr2(Deng et al., 2022)
$Venue$ ICCV’23 MM’23 AAAI’23 ICLR’22 - CVPR’23 ICCV’23 CVPR’22
$LP_{\text{bg}}\downarrow$ 0.11/0.12 0.20/0.36 0.12/0.12 0.37 0.24 0.72/0.56 0.69 0.71/0.71 0.61
$LP_{\text{fg}}\downarrow$ 0.10/0.32 0.32/0.36 0.10/0.39 0.41 0.21 0.11/0.45 0.51 0.11/0.32 0.40
$CP_{\text{img}}\uparrow$ 95.42/78.63 85.35/82.05 95.13/73.65 78.85 87.55 96.43/69.57 80.27 95.54/73.91 83.57
$CP_{\text{style}}\uparrow$ 47.50/56.37 47.66/47.40 47.64/55.96 50.62 49.28 57.97/78.60 64.59 59.25/67.04 63.28
$CP_{\text{dir}}\uparrow$ 0.11/11.69 2.96/4.63 0.35/15.39 8.50 3.27 3.97/51.59 26.90 6.22/35.63 22.08
5. Experiments
5.1. Setup
@@
We employ the Stable Diffusion model (Rombach et al., 2022) as the pretrained backbone and utilize DPM Solver++ as the scheduler with a total of 25 steps for both inversion and reconstruction. Specifically, we first resize the input images $I^{\text{f}}$, $I^{\text{b}}$, and $I^{\text{comp}}$ to 512$\times$512, and encode them into corresponding $z^{\text{f}}_{0}$, $z^{\text{b}}_{0}$, and $z^{\text{comp}}_{0}$. Afterward, we take these latents with prompt embedding $\rho_{exceptional}$ as the input of Stable Diffusion model during both inversion and reconstruction stage. For the style transfer task, we set $\alpha$ to 0.9, $\beta$ to 1.1, $T_{share}$ to 25 (we activate the share-attention layer when t¡$T_{share}$), and $L_{share}$ to 14 (out of 16 total layers in the diffusion U-Net). As for the object swapping purpose, we change $T_{share}$ to 20, and for the object insertion usage, we change $T_{share}$ to 15. We refer to these hyperparameters ($T_{share}$, $L_{share}$, $\alpha$, $\beta$) as “inference-time-adjustable hyperparameters” since they can be flexibly adjusted to modulate the strength of style according to different use cases during the inference process. Due to the space constraint, please refer to Appendix  G.1 in the supplementary materials for the sensitivity test on the hyperparameters.
5.2. Datasets and Metrics
@@
Datasets. We generalize the computational metrics and benchmarks from various image editing methods: “Painterly image harmonization” (Lu et al., 2023a; Cao et al., 2023a; Zhang et al., 2020), “Prompt-based image composition” (Lu et al., 2023b; Meng et al., 2022; Gu et al., 2024), “Style Transfer” (Huang et al., 2023; Cheng et al., 2023; Zhang et al., 2023; Deng et al., 2022), and our proposed “General Painterly Harmonization” achieved by the General Painterly Harmonization Benchmark (GPH Benchmark). This benchmark generalizes real-world usage scenarios of the aforementioned methods including “Object Insertion”, “Object Swapping” and “Style Transfer” providing a more practical benchmark for evaluating the applicability of cross-image editing methods and aims to mitigate the shortcomings of existing benchmarks such as WikiArt combined with COCO (Tan et al., 2019; Lin et al., 2014) and the TF-ICON Benchmark (Lu et al., 2023b). Details of these datasets and the motivation of GPH Benchmark can be found in Appendix  B.
@@
LPIPS and CLIP regarding computation metrics. LPIPS (Zhang et al., 2018) and CLIP (Radford et al., 2021) (abbreviated as $LP$ and $CP$ respectively in the subsequent discussion) are two widely utilized metrics for evaluating image similarity, each leveraging unique aspects of image analysis. LPIPS employs a pretrained model from ImageNet to extract deep features, assessing perceptual similarity based on these features. In contrast, CLIP, which is trained on a large-scale dataset of images and text, excels at capturing high-level semantic information in both visual and contextual domain. In TF-ICON benchmark (Lu et al., 2023b), these two metrics are leveraged to assess content preservation and stylized performance, where $LP_{\text{fg}}$ and $CP_{\text{img}}$ are calculated to measure the content consistency and image semantic similarity, respectively. Moreover, $LP_{\text{bg}}$ is also used to measure background consistency before and after harmonization. On the other hand, StyleGAN-NADA (Gal et al., 2022) proposes $CP_{\text{dir}}$ to calculate the alignment level between the feature shift direction of pasted object and the background. Specifically, let $f(\cdot)$ denote the extracted feature of CLIP image encoder. The alignment level is calculated by the similarity between $\{f(I^{o})-f(I^{f})\}$ and $\{f(I^{b})-f(I^{f})\}$. Finally, following VCT (Cheng et al., 2023), we adopt $CP_{\text{style}}$ to measure the feature similarity of harmonized images and style references.
@@
Range-based evaluation. While metrics such as LPIPS and CLIP are useful for assessing content fidelity and stylization in image harmonization, they can sometimes emphasize either too much content preservation or excessive stylization, as shown in Fig. 11. This imbalance may compromise the overall harmonization and visual quality of the output image. To address this issue, we propose that an effective general painterly harmonization pipeline should offer users the flexibility of balancing between stylization intensity and content integrity by adjusting hyperparameters. To quantitatively measure this capability, we suggest defining upper and lower bounds for content preservation and stylization, which can serve as indicators of a method’s adaptability across different harmonization scenarios. In other words, methods with a wider range should be able to adapt to more use cases. Besides, we have re-implemented existing baselines with configurable hyperparameters to facilitate a range-based comparison.  [^2] More implementation details can be found in Appendix  C.
Figure 5. Comparison of different stylized strength.
5.3. Qualitative Comparison
@@
For qualitative comparison, TF-GPH showcases remarkable capabilities in our proposed GPH Benchmark as depicted in Fig. 4, ranging from low-level texture harmonization like transitioning to singular colors and color matching (rows 1 and 2) to high-level semantic harmonization such as extending the skin color of the replaced man onto the pasted face or redrawing the covered beard along with the chin line of the pasted face (rows 3 and 4). While PHDiff and SDEdit are able to achieve low-level texture harmonization, they often struggle with high-level semantic blending, failing to successfully recover face feature. Additionally, our method is also able to seamlessly composites images without the need for prompts in TF-ICON Benchmark(see Appendix  D.2) and showing outstanding performance in other existing dataset (see Appendix  D.1). This highlights how our similarity reweighting technique effectively leverages the innate characteristics of the diffusion model to achieve both texture and semantic harmonization with image-wise attention.
Moreover, our proposed TF-GPH demonstrates exceptional performance in style transfer (row 5,6,7 in Fig. 4), utilizing a diverse range of graphic art forms within the test dataset. Our method excels at stylizing original content while maintaining superior image quality, by leveraging advanced diffusion image generation techniques. This addresses a notable limitation observed in non-diffusion methods such as QuantArt and StyTr2, which often struggle to maintain image quality. Additionally, TF-GPH better preserves the layout of the content image compared to tuning-based diffusion editing methods like VCT, credits to its innovative share-attention layer. This layer effectively retains feature information from reference images during inference, thereby enhancing content detail in the generated output. More qualitative results are available in Appendix  D.
5.4. Quantitative results
Tab. 1 and Tab. 2 showcase the quantitative results from our proposed GPH and TF-ICON Benchmark respectively. Notably, a detailed range-based evaluation on WikiArt combined with COCO is documented in Appendix  E.1. TF-GPH consistently demonstrates superior performance across these benchmarks, outperforming existing evaluation criteria. Our uniquely designed share-attention module significantly enhances reference information retention compared to prompt-based editing methods like TF-ICON and SDEdit, improving $LP_{\text{bg}}$, $LP_{\text{fg}}$ lower bounds by $45\%$ and $52\%$ at least . These methods often struggle to preserve details of the composite image when the prompts inadequately complement the reference images.
Furthermore, TF-GPH utilizes a pre-trained diffusion model augmented with a similarity reweighting mechanism, achieving higher stylization upper bounds than methods that require extensive training, such as QuantArt and StyTr2. These conventional methods frequently encounter limitations in handling diverse styles within the test data, while our TF-GPH achieves a substantial improvement around $17\%\sim 28\%$ in $CP_{\text{style}}$ and $44\%\sim 133\%$ in $CP_{\text{dir}}$ upper bounds compared to other style transfer methods. Importantly, TF-GPH’s inference-time-adjustable hyperparameters enable it to offer the broadest range of content preservation and stylization capabilities, as evidenced in both Tab. 1 and Appendix  E.1.
@@ 5.4. Quantitative results
We also conduct a comprehensive user study to evaluate the efficacy of TF-GPH. The study encompasses two distinct evaluation tasks: Style Transfer and Painterly Harmonization, which includes object insertion and swapping. For this, we recruited 50 participants through Amazon MTurk, each tasked with responding to 10 questions. Participants were instructed to assess the generated images based on three criteria: (1) Content Consistency, (2) Style Similarity, and (3) Visual Quality, rating them on a scale from 1 (worst) to 4 (best). Specifically, for the Painterly Harmonization task, ’Content Consistency’ was subdivided into ’Object Consistency’ and ’Background Consistency’.
Fig. 6(a) displays the results for the painterly harmonization task, where TF-GPH not only excels in stylization but also in preserving superior visual quality of the pasted objects. Additionally, Fig. 6(b) illustrates the outcomes for the style transfer task, with TF-GPH achieving the highest scores in overall quality and content consistency, along with competitive style similarity. These results validate our hypothesis that visual quality transcends mere content preservation or style strength. The consistently high performance across these metrics underlines the robustness of our method for diverse real-world applications. Further details on the methodology and results of the user study are provided in Appendix  F.
5.5. Ablation Study
@@
Tab. 3 reveals the impact of components within TF-GPH. Simply applying reconstruction to the composite image is ineffective at harmonizing pasted object into background. Additionally, incorporating a vanilla share-attention layer offers minimal improvement due to the inherent nature of the attention mechanism, as previously discussed. However, the integration of the similarity reweighting strategy within the share-attention layer allows the TF-GPH pipeline to successfully encode cross-image information into the composite image, resulting in significant enhancements in the stylization indexes $CP_{\text{style}}$ and $CP_{\text{dir}}$ across both tasks.
Figure 6. User study result, we abbreviate “Visual Quality”, “Style Similarity”, “Object Consistency”, “Background Consistency” and “Content Consistency” into “VQ”, “SS”, “OC”, “BC”, and “CC”, respectively.
Table 2. Quantitative result of TF-ICON Benchmark
SDEdit (0.4)(Meng et al., 2022) 0.35 0.62 80.56 27.73
Blended(Avrahami et al., 2023) 0.11 0.77 73.25 25.19
Paint(Yang et al., 2023) 0.13 0.73 80.26 25.92
DIB(Zhang et al., 2020) 0.11 0.63 77.57 26.84
TF-ICON(Lu et al., 2023b) 0.10 0.60 82.86 28.11
Ours 0.05 0.48 83.34 30.33
@@
Table 3. Ablation study on TF-GPH’s components in painterly harmonization (upper) and style transfer (bottom) on GPH. (We abbreviate “Share-Attention layer” and “Similarity Reweighting” as “SA” and “SR” )
Metrics $LP_{\text{bg}}\downarrow$ $LP_{\text{fg}}\downarrow$ $CP_{\text{style}}\uparrow$ $CP_{\text{dir}}\uparrow$
Vanilla 0.11 0.09 47.50 0.11
+$SA$ 0.11/0.11 0.10/0.10 47.50/47.50 0.11/0.18
+$SR$ 0.11/0.12 0.10/0.32 47.50/56.37 0.11/11.69
Vanilla 0.72 0.11 57.97 3.97
+$SA$ 0.72/0.69 0.11/0.12 57.97/59.05 3.97/5.12
+$SR$ 0.72/0.56 0.11/0.45 57.97/78.60 3.97/51.59
6. Conclusion

### v2

@@ Introduction
where $\odot$ denotes the Hadamard product. Afterward, Eq.(2) outputs the batch attention $\hat{A}\in\mathbb{R}^{(3\cdot m)\times d}$ containing the intact $A^{\text{f}}$, $A^{\text{b}}$, and $A^{\text{c}}$ guided by the features of $z^{\text{b}}$ and $z^{\text{b}}$. The specially designed $\hat{M}$ can be visualized as:
@@
Here, $J\in\mathbb{R}^{m\times m}$ is an all-one matrix, and $\nu=-\infty$ minimizes the similarity between $Q$ and $K$ on the corresponding entry, keeping $A^{\text{f}}$, $A^{\text{b}}$ intact. While $\alpha$, $\beta$, and $\gamma$ control the attention of $Q^{\text{c}}$ towards $K^{\text{f}}$, $K^{\text{b}}$, and $K^{\text{c}}$, respectively.. It is worth noting that when setting $\alpha=-\infty$, $\beta=-\infty$, and $\gamma=1$, each row in Eq.(2) is equivalent to Eq. (1) as $Q^{\text{c}}$, $Q^{\text{f}}$, and $Q^{\text{b}}$ can only attend to its counterpart $K^{\text{c}}$, $K^{\text{f}}$, and $K^{b}$ without information from other images. Therefore, our proposed similarity disentangle mask can be viewed as an expansion of attention mechanism with adjustable entries controlling features sharing.
Furthermore, to completely disentangle the features related to the object reference $z^{\text{f}}$ from $z^{\text{b}}$, we set the entry $\gamma$ to $\infty$- ∞, which blocks the functionality of $K^{\text{c}}$ and $V^{\text{c}}$. By this means, we can control the features related to the pasted-foreground object within $z^{\text{c}}$ by modifying entry $\alpha$, which control the influence of $z^{\text{f}}$, and similarly, control the background features within $z^{\text{c}}$ related to $z^{\text{b}}$ by adjusting its corresponding entry $\beta$. As shown in Fig. 4(b), the output remains nearly the same to Fig. 4(a) validating that the features of $z^{\text{c}}$ can be totally controlled by two references $z^{\text{f}}$ and $z^{\text{b}}$
@@
Figure 4: Comparisons of different attention strategy with corresponding similarity mask (read with Fig. 3).
Similarity Reweighting
@@
Another intriguing observation from Fig. 4(a) and (b) is that the output image only changes slightly even when the pasted “corgi” in $I^{\text{c}}$ has a different resolution compared to the “corgi” in $I^{f}$. We infer that the robustness of the pretrained diffusion model enables it to capture high-level semantic and structural information despite minor disturbances, such as differences in scale and position. Consequently, the self-attention layer can withstand these perturbations, producing results that remain similar to the original input.
To determine the perturbation that can break the self-attention robustness while still generating high-quality results, a simple yet effective idea of attention injection has been widely adopted by previous research (Gu et al. 2024; Hertz et al. 2023; Lu, Liu, and Kong 2023; Tumanyan et al. 2023). This approach introduces strong perturbations to the attention mechanism by directly appending either prompt-guided cross-attention output or image-guided self-attention output computed with other images. Another common strategy is applying the “AdaIN” technique to different components. For example, (Chung, Hyun, and Heo 2024) compute the mean and variance of $z^{\text{b}}$, then normalize $z^{\text{c}}$ with these computed values, or (Hertz et al. 2024) perform AdaIN normalization on $K^{\text{b}}$ and $K^{\text{c}}$. Although these direct modifications to $z^{\text{c}}$ can produce exaggerated image editing effects, they also disrupt semantic details and structural coherence.
In contrast to the aforementioned strategies, we argue that certain attributes crucial to content identity should not be entirely replaced by style features, as discussed in (Saini, Pham, and Shrivastava 2022). For example, the yellow hue of a corgi is an essential part of its identity and should be preserved rather than changed to the global background color tone such as blue or black. Instead, integrating the yellow color from the background style into the corgi would better maintain its content identity as shown in Fig. 4(c). To achieve this, we prioritize high-similarity style features that potentially possess content-related attributes such as color, texture, or semantics. Instead of evenly scaling the attention output as in (Deng et al. 2023), scaling the similarity has a different effect due to the softmax process involved. By scaling the input similarity before applying softmax, high-similarity features are amplified while low-similarity features are diminished in the final attention output. This approach helps minimize content disruption during stylization. Without loss of generality, we place a higher tendency on style reference $z^{\text{b}}$ by setting $\beta$ to 1.1, and a minor preference on content preservation related to $z^{\text{f}}$ by setting $\alpha$ to 0.9, our designed TF-GPH achieves remarkable painterly harmonization effects without losing content structure and background consistency. The overall algorithm and visualization can be found in Appendix.
@@
Figure 5: Qualitative result of object insertion (rows 1 and 2), object swapping (rows 3 and 4), and style transfer (rows 5 and 6)
@@
Table 1: Quantitative results of GPH-Benchmark ( represents the method with inference-time-adjustable hyperparameters. The left side of / represents content emphasis strategy, while the right side of / represents stylized emphasis strategy.)
Painterly Harmonization (512x512) Style Transfer (512x512)
Ours ArtoPIH ProPIH TF-ICON PHDiff Ours StyleID Z-STAR StyTr2
$Venue$ AAAI’24 AAAI’24 ICCV’23 MM’23 - CVPR’24 CVPR’24 CVPR’22
$LP_{\text{bg}}\downarrow$ 0.11/0.12 0.25 0.31/0.31 0.20/0.36 0.12/0.12 0.72/0.55 0.60/0.56 0.70/0.63 0.61
$LP_{\text{fg}}\downarrow$ 0.10/0.32 0.37 0.34/0.42 0.32/0.36 0.10/0.39 0.11/0.45 0.36/0.48 0.15/0.37 0.40
$CP_{\text{img}}\uparrow$ 95.42/78.63 84.96 87.78/77.24 85.35/82.05 95.13/73.65 96.43/69.57 83.26/69.80 92.31/77.20 83.57
$CP_{\text{st}}\uparrow$ 47.50/56.37 49.67 47.87/51.19 47.66/47.40 47.64/55.96 57.97/78.60 67.70/77.47 59.61/69.50 63.28
$CP_{\text{dir}}\uparrow$ 0.11/11.69 5.40 2.83/10.09 2.96/4.63 0.35/15.39 3.97/51.59 26.96/50.24 9.76/34.83 22.08
Experiments
@@
Setup. We employ the Stable Diffusion model (Rombach et al. 2022) as the pretrained backbone and utilize DPM Solver++ as the scheduler with a total of 25 steps for both inversion and reconstruction. Specifically, we first resize the input images $I^{\text{f}}$, $I^{\text{b}}$, and $I^{\text{c}}$ to 512$\times$512, and encode them into corresponding $z^{\text{f}}_{0}$, $z^{\text{b}}_{0}$, and $z^{\text{c}}_{0}$. Afterward, we take these latents with prompt embedding $\rho_{exceptional}$ as the input during both inversion and reconstruction stage. As for the rest of setting, we refer these hyperparameters ($T_{share}$, $L_{share}$, $\alpha$, $\beta$) as “inference-time-adjustable hyperparameters” since they can be flexibly adjusted to modulate the strength of style according to different use cases during the inference process, we leaves remain setting details in the Appendix.
@@
Datasets. We generalize the computational metrics and benchmarks from various image editing methods including “Painterly image harmonization”, “Prompt-based Image Composition”, and “Style Transfer”. Additionally, we examined different approaches on our proposed “General Painterly Harmonization” achieved by the General Painterly Harmonization Benchmark (GPH Benchmark). This benchmark generalizes real-world usage scenarios of the aforementioned methods including “Object Insertion”, “Object Swapping” and “Style Transfer”, providing a more practical benchmark and aims to mitigate the shortcomings of existing benchmarks such as WikiArt combined with COCO (Tan et al. 2019; Lin et al. 2014) and the TF-ICON Benchmark (Lu, Liu, and Kong 2023). Details and experiment of these datasets can be found in the Appendix.
@@
LPIPS and CLIP regarding computation metrics. In our evaluation, we use LPIPS (Zhang et al. 2018) and CLIP (Radford et al. 2021) metrics, abbreviated as $LP$ and $CP$ respectively. LPIPS is sensitive to low-level visual features, while CLIP excels in capturing high-level semantic features. In TF-ICON benchmark, these two metrics are leveraged to assess content preservation and stylized performance, where $LP_{\text{fg}}$ and $CP_{\text{img}}$ are calculated to measure the content consistency and image semantic similarity, respectively. Moreover, $LP_{\text{bg}}$ is also used to measure background consistency before and after harmonization. And $CP_{\text{dir}}$ (Gal et al. 2022) to calculate the alignment level between the feature shift direction of pasted object and the background. Finally, we adopt $CP_{\text{st}}$ (Cheng et al. 2023) to measure the feature similarity of harmonized images and style references.
@@
Range-based evaluation. While metrics such as LPIPS and CLIP are useful for assessing content fidelity and stylization in image harmonization, they can sometimes emphasize either too much content preservation or excessive stylization, resulting inharmonious image. Therefore, we argue that an effective pipeline should offer users the flexibility of balancing between stylization intensity and content integrity by adjusting hyperparameters. To measure this capability, we suggest defining upper and lower bounds for content preservation and stylization, which can serve as indicators of a method’s adaptability across different harmonization scenarios. The corresponding upper and lower settings for the baselines are provided in the Appendix.
Baselines. We compared our proposed TF-GPH with different state-of-the-art methods on various tasks for the comprehensive assessment. For “Painterly Image Harmonization”, we incorporate ArtoPIH (Niu et al. 2024a), ProPIH (Niu et al. 2024b), PHDiffusion (Lu et al. 2023), and PHDNet (Cao, Hong, and Niu 2023). For “Prompt-Based Image Composition”, we use TF-ICON (Lu, Liu, and Kong 2023). In the “Style Transfer” category, we evaluate Style Injection (StyleID) (Chung, Hyun, and Heo 2024), ZSTAR (Deng et al. 2023), StyTr2 (Deng et al. 2022), QuantArt (Huang et al. 2023), and VCT (Cheng et al. 2023). For models designed for 256x256 resolutions, e.g. ProPIH, we resized the output to 512x512 for high-resolution evaluation. Comparisons of 256x256 resolution are in the Appendix.
Qualitative Comparison
@@
For qualitative comparison, TF-GPH showcases remarkable capabilities in our proposed GPH Benchmark as depicted in Fig. 5, ranging from low-level texture harmonization such as transitioning to singular colors and color matching (rows 1 and 2) to high-level semantic harmonization such as extending the skin color of the replaced man onto the pasted face or redrawing the covered beard along with the chin line of the pasted face (rows 3 and 4). Although ArtoPIH and ProPIH are able to achieve low-level texture harmonization, they struggle with high-level semantic blending, such as face recovery in the row 3 of Fig. 5, due to training data limitation. This highlights how our similarity reweighting technique effectively leverages the characteristics of the diffusion model to achieve both texture and semantic harmonization with image-wise attention.
Moreover, our proposed TF-GPH demonstrates exceptional performance in style transfer (row 5,6 in Fig. 5). Our method outperforms others in stylizing original content while maintaining high image quality, effectively mitigating the common issue of content disruption seen in other attention-based methods. For instance, our approach preserves content coherence more accurately than StyleID and ZSTAR as shown in row 5. Furthermore, our model excels in blending photographic features into sketches, where other methods fail, as depicted in the row 6. This illustrates that our similarity disentangle mask not only preserves content information effectively but also extracts style features robustly, even in scenarios like photography.
Quantitative Results
Tab. 6 presents the quantitative results of the GPH Benchmark. The performance of TF-GPH consistently surpasses that of existing evaluation criteria on different benchmarks. Our similarity disentangle mask significantly improves reference preservation compared to prompt-based editing methods such as TF-ICON, as well as traditional harmonization methods like ArtoPIH and ProPIH, achieving the lowest $LP_{\text{bg}}$ and $LP_{\text{fg}}$ values while also demonstrating superior stylization with the highest $CP_{\text{st}}$.
Moreover, TF-GPH employs a novel similarity-based editing technique that consistently outperforms existing attention-based methods, such as StyleID, in both content preservation metrics ($LP_{\text{fg}}$, $CP_{\text{img}}$) and stylization metrics ($CP_{\text{st}}$, $CP_{\text{dir}}$). Additionally, the wide content preservation and stylization range of TF-GPH confirm the potential of our inference-time-adjustable hyperparameters, which can accommodate various preferences.
@@
We also conduct user preference studies, which are viewed more reliable (Podell et al. 2023). The study encompasses two tasks: Style Transfer and Painterly Harmonization, which includes object insertion and swapping. For each task, we recruited 20 participants, each asked with responding to 20 image pairs. Participants were instructed to compared the generated images based on three criteria: (1) Content Consistency, (2) Style Similarity, and (3) Visual Quality. We provide the results in Fig. 6, where TF-GPH achieving the highest preference in overall quality and content consistency, along with competitive style similarity. These results validate our hypothesis that visual quality transcends mere content preservation or style strength.
Ablation Study
@@
Tab. 2 reveals the impact of components within TF-GPH. Simply applying reconstruction to the composite image is ineffective at harmonizing pasted object into background. By contrast, when we incorporating similarity disentangle mask, we perfectly disentangle the attention of $z^{\text{c}}$ to the two other image latents $z^{\text{f}}$, $z^{\text{b}}$ and reach nearly no reconstruction loss. Furthermore, the integration of the similarity reweighting strategy significantly improves the stylization indices $CP_{\text{st}}$ and $CP_{\text{dir}}$ across both tasks, demonstrating its effectiveness in encoding cross-image information into the composite image.
Figure 6: User preferencestudy result.
@@
Table 2: Ablation study on TF-GPH’s components in painterly harmonization (upper) and style transfer (bottom) on GPH Benchmark. (We abbreviate “Similarity Disentangle Mask” and “Similarity Reweighting” as “SDM” and “SR” ). Because the +SDM+SR is the stylization emphasis strategy of +SDM, we put its stylization upper bound here.
Metrics $LP_{\text{bg}}\downarrow$ $LP_{\text{fg}}\downarrow$ $CP_{\text{st}}\uparrow$ $CP_{\text{dir}}\uparrow$
Reconstruction 0.11 0.09 47.50 0.11
+$SDM$ 0.11 0.10 47.50 0.18
+$SDM$+$SR$ /0.12 /0.32 /56.37 /11.69
Reconstruction 0.72 0.11 57.97 3.97
+$SDM$ 0.69 0.12 59.05 5.12
+$SDM$+$SR$ /0.56 /0.45 /78.60 /51.59
Conclusion